# OG-VLA: 3D-Aware Vision Language Action Model via Orthographic Image Generation

### v1

@@ C.3 Model Ablations
C.4 Qualitative results
C.5 Ablation with ground truth translation and rotation
@@
Table 7: Ablated evaluation of the model’s translation and rotation prediction capabilities on a sampled training set of 20 episodes, similar to the test sets. In each evaluation, we individually ablate either the translation or rotation predictions to assess the model’s prediction capabilities for each.
We present another set of ablations in Table 7 for studying our model’s translation and rotation prediction abilities in an ablated setting on a sampled training set. First, we note that the overall performance of our model on Training and Novel Pose split is quite similar, which indicates that the model has not overfit and is generalizing well w.r.t the training performance. We observe that there is huge scope for improvement, both in predicting more precise translations and rotations from the last two rows of the table. We believe that this gap can be closed with training techniques like co-training with other datasets for better reasoning [42] and leveraging existing large robotics datasets like that in other VLA models [3]. We believe leveraging these datasets with a 3D-aware VLA framework can unlock better learning signals from these large-scale datasets and significantly improve results for our proposed VLA architecture.
D Real world details

### v2

@@ C-B Image Generation Mode
Table IX: Ablated evaluation of the model’s translation and rotation prediction capabilities on a sampled training set of 20 episodes, similar to the test sets. In each evaluation, we individually ablate either the translation or rotation predictions to assess the model’s prediction capabilities for each.
We report success rates across tasks and test splits in Table VII. Generation without reconstruction of the scene performs the worst of all generation modes. This may be due to forcing the ImageGenerator to unlearn it’s prior of generating natural images and forcing it to only predict Gaussian distributions. We also observe instability in training this version, as sometimes the ImageGenerator would collapse to produce completely black or noisy images and stop generating actions on them. Generation with reconstruction and that with faded reconstruction did not show consistent difference over all test splits. Generation with reconstruction learns a policy that’s better by 4.4% on Novel Pose split and 2.3% Novel State split. Generation with faded reconstruction performs better on Novel Object and Novel State splits by 3.3% and 3%. Therefore, we conclude that these model variants have similar performance, and generation with faded reconstruction did not work better as we had hypothesized due to its balance between forcing the model to focus less on scene reconstruction and more on action prediction. Therefore, we report generation with reconstruction as our final method.
@@
C-C Model Ablations
We present ablations on OG-VLA’s design choices in Table VIII. The first result shows the effect of tiling the 4 orthographic views instead of feeding them to the ImageGenerator in batch of 4 as also explored in Genima [shridhar2024generative]. For tiling, we stack the 4 views as in 2D array, following prior work. Tiling did not improve the performance for our model as opposed to results reported in the prior works. This may have occurred because the prior work did not have an LLM in the pipeline, so tiling becomes necessary for modeling interactions between views to generate multi-view consistent predictions. However, due to the LLM in OG-VLAconditioning generation in all views, there’s sufficient interaction and reasoning between views through the predicted image tokens $\left(t^{i}_{a},\quad i\in\{1,\dots,4\}\right)$.
Finally, we study the effect of bypassing the prompt and instruction into the ImageGenerator, shown in the last section of Table VIII. The performance drop suggests that some instruction information may have been lost in image tokens output from the LLM, therefore it is beneficial to provide that information separately to the ImageGenerator for more accurate action prediction.
C-D Qualitative results
C-E Ablation with ground truth translation and rotation
We present another set of ablations in Table IX for studying our model’s translation and rotation prediction abilities in an ablated setting on a sampled training set. First, we note that the overall performance of our model on Training and Novel Pose split is quite similar, which indicates that the model has not overfit and is generalizing well w.r.t the training performance. We observe that there is huge scope for improvement, both in predicting more precise translations and rotations from the last two rows of the table. We believe that this gap can be closed with training techniques like co-training with other datasets for better reasoning [yuan2024robopoint] and leveraging existing large robotics datasets like that in other VLA models [kim24openvla]. We believe leveraging these datasets with a 3D-aware VLA framework can unlock better learning signals from these large-scale datasets and significantly improve results for our proposed VLA architecture.
D Real world details